I stand on a building 640 feet high and throw a rock into the air with a velocity of 96 ft/sec. When does the rock have a height of 800 feet?
Answer by stanbon(75887)

I stand on a building 640 feet high and throw a rock into the air with a velocity of 96 ft/sec. When does the rock have a height of 800 feet?
---
h(t) = -16t^2 + 96t + 640
------
Solve:
-16t^2 + 96t + 640 = 800
---
-16t^2 + 96t - 160 = 0
----
-16(t^2 - 6t + 10) = 0
---
Quadratic formula:
t = [6 +- sqrt(36-4*10)]/(2)
----
Since 36-4*10 is negative there is no Real Number solution.
---
The rock never teaches a height of 800 ft.
=============
Cheers,
Stan H.
==============